OpenStudy (anonymous):

simplify 15-2^3/4*7- 7

OpenStudy (callisto):

Is this your question? \[\frac{15-2^3}{4*7- 7}\]

OpenStudy (anonymous):

yes how did you type it this way meaning that line in between

OpenStudy (callisto):

\frac{15-2^3}{4*7- 7} ^ numerator ^ donminator the add \[ at the front of \frac

OpenStudy (callisto):

and add \] at the end

OpenStudy (callisto):

\[\frac{15-2^3}{4*7- 7}= \frac{15-8}{28- 7} = \frac{7}{21} =???\]

OpenStudy (anonymous):

3

OpenStudy (anonymous):

thanks for showing me how to do this I have gotten alot of answer wrong in my class for not setting the problem up correctly.

OpenStudy (callisto):

Hm... it's not 3!!!

OpenStudy (callisto):

\[\frac{7}{21} = \frac{7 \times 1}{7 \times 3}\] Cross out the common factor, what do you get?

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

{1/3}

OpenStudy (callisto):

Yes!~